OpenStudy (anonymous):

The equation of the line that goes through the points (5,9) and (6,3) can be written in the form y=mx+b, where is m? where is b?

OpenStudy (campbell_st):

find the gradient (slope) which is rise/run \[m = (y _{2} - y _{1})/(x _{2} -x _{1})\] let (5,9) be the 2 point (6, 3) be the 1 point m = (9 - 3)/(5 - 6) m = -6 use the point slope formula \[y - y _{1} = m(x - x _{1})\] y - 3 = -6(x - 6) expand and simplify y = -6x + 39

OpenStudy (campbell_st):

m is the slope of the line, b is the y-intercept, where x = 0 the point is (0, 39)